A car that is traveling down a highway has
1,30 x 10s J of kinetic energy. A truck with 8.30
times the mass of the car is traveling at the same
speed. What is the kinetic energy of the truck
1
Expert's answer
2020-12-15T11:46:31-0500

Let's denote "E =1.3 \times10 \,J = 13 J"

"m_1" - mass of the car, "m_2" - mass of the truck. Also, "\frac{m_2}{m_1} =8.3" and "v_1 = v_2".

The kinetic energy of the body is

"\displaystyle E = \frac{mv^2}{2}"

"\displaystyle \frac{E_2}{E_1} = \frac{m_2v_2^2}{m_1v_1^2} = \frac{m_2}{m_1} =8.3"

So, "E_2 = 8.3 E_1 = 8.3 \cdot 13 = 107.9 \, J"

Answer: 107.9 J